OpenStudy (anonymous):

how do you factor 2u n-1=-0.85u n-1?

jhonyy9 (jhonyy9):

2u(n-1)=-0,85u(n-1) 2u(n-1)+0,85u(n-1)=u(n-1)(2+0,85)

jhonyy9 (jhonyy9):

2,85u(n-1)